OpenStudy (anonymous):

I need someone to explain how to solve this equation!!! 0.8+6 over 4 (1/5+5)

OpenStudy (aric200):

Is this how it is written out or did you input it this way?

OpenStudy (anonymous):

I just inputed it that way. The original problem is b+6 (over/fraction bar) 4 (d+c) which then becomes 0.8+6 (over/fraction bar) 4 (1/5 + 5)

OpenStudy (zzr0ck3r):

\[\dfrac{0.8+6}{4(\frac{1}{5}+5)}=\frac{\frac{8}{10}+6}{\frac{4}{5}+20}=\dfrac{\frac{8}{10}+\frac{60}{10}}{\frac{4}{5}+\frac{100}{5}}=\dfrac{\frac{68}{10}}{\frac{104}{5}}=\dfrac{34}{5}\dfrac{5}{104}=\dfrac{34}{104}=\dfrac{17}{52}\]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (zzr0ck3r):

np, let me know if you have any questions about the steps